OpenStudy (anonymous):

5+i and 5 - i find a polynomial function f(x) of least degree having only real coefficients with zeros.

hartnn (hartnn):

if 'a' is the root, then (x-a) is the factor! if 5+i is the root, then (x-(5+i)) = (x-5-i) is the factor, if 5-i, is the root, then ???

OpenStudy (anonymous):

(x-(5-i)) = (x+5+i) ???

hartnn (hartnn):

(x-(5-i)) = (x-5+i)

hartnn (hartnn):

so you got the 2 factors, now just multiply them :) f(x) = (x-5-i)(x-5+i) multiply those 2 factors out

OpenStudy (anonymous):

thanks so much for the help you can't imagine how lost i was

hartnn (hartnn):

glad to help :) welcome ^_^